Sumter Company manufactures televisions and they had a very successful year. They had 225,000 shares of common stock outstanding and they have decided to do a 3-for -1 stock split. Instructions: a) What will be the number of shares outstanding after the split? b) If the common stock had a par value $30 per share be fore the stock split, what would be the par value after the stock split? c) If the common stock had a market price of $150 per share before the stock split, what would be the approximate market price after the stock split?
Answer #1

a) Outstanding shares after stock split = 225000*3 = 675000 Shares

b) Par value after stock split = 30/3 = $10 per share

c) Market price after stock split = 150/3 = $50 per share